# Lower semicontinuity, Stoilow factorization and principal maps

### v2

@@ 3.2. Lower semicontinuity with integrable distortion
as $R\nearrow 1$, Inserting the Jensen inequality for each $\nu_{z}$ in (3.13) yields the required lower semicontinuity. ∎
@@ 3.3.1. An integrability property of the Burkholder functional
[…]
@@
Changing variables we may without loss of generality assume that $z_{0}=0$ and $R=1$. We then Stoilow factorize $u$ on $\mathbb{D}$, whereby $u=h\circ f$ for a $K$-quasiconformal principal map $f\colon\mathbb{C}\to\mathbb{C}$ and a holomorphic map $h\colon f(\mathbb{D})\to\mathbb{C}$. If $m\equiv\sup_{\mathbb{D}}|u|$, then also $m=\sup_{f(\mathbb{D})}|h|$ and so by Cauchy’s integral formula and Hölder properties of $K$-quasiconformal principal maps [5] we estimate
@@
We split the proof into two steps (building on our prerequisites) and remark that the assumption $u\in\mathrm{W}^{1,p}_{\operatorname{loc}}(\Omega)$ only is used in the second step.
By Proposition 3.6, the sequence $\bigl{(}\mathbf{B}_{p}(\textup{D}u_{j})\bigr{)}$ is bounded in $\mathrm{L}^{1}_{\operatorname{loc}}(\Omega)$. Because $\mathrm{L}^{1}_{\operatorname{loc}}(\Omega)\subset\mathrm{C}_{c}(\Omega)^{\ast}$ continuously, the Banach-Alaoglu compactness theorem implies that any subsequence of $\bigl{(}\mathbf{B}_{p}(\textup{D}u_{j})\bigr{)}$ admits a further subsequence that converges $\mbox{}^{\ast}$ to a Radon measure on $\Omega$. Any such limit must necessarily be a nonpositive Radon measure. Assume that $-\lambda$ is such a $\mbox{}^{\ast}$ limit and consider the Lebesgue-Radon-Nikodym decomposition
@@
Step 1. $-\frac{\mathrm{d}\lambda}{\mathrm{d}\mathscr{L}^{2}}\geq\mathbf{B}_{p}(\textup{D}u)$ holds $\mathscr{L}^{2}$ almost everywhere in $\Omega$. The proof consist of a rather general localization argument at a point of differentiabilty of $u$. Arguments of this sort are known in the literature as blow-up arguments and in the present context go back to I. Fonseca and S. Müller [25]. The fact that the Burkholder functional is nonpositive on its effective domain means that one must be more careful than in the classical set-up, where the considered functionals are bounded below and therefore concentration effects are not so harmful.
@@
(3.16) $u\mbox{ is differentiable at }z_{0}\mbox{ with Jacobi matrix }A_{0}\equiv\textup{D}u(z_{0})$
@@
$\lim_{E\notowner r\searrow 0}\lim_{j\to\infty}\mathchoice{{\vbox{\hbox{$\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\scriptstyle-$}}\kern-3.25pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.29166pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-1.875pt}}\!\int_{\mathbb{D}_{r}(z_{0})}\!\mathbf{B}_{p}(\textup{D}u_{j})\,\mathrm{d}m(z)=-\frac{\mathrm{d}\lambda}{\mathrm{d}\mathscr{L}^{2}}(z_{0}),$
@@
Finally, taking $s\nearrow 1$ we conclude that $-\frac{\mathrm{d}\lambda}{\mathrm{d}\mathscr{L}^{2}}(z_{0})\geq\mathbf{B}_{p}(\textup{D}u(z_{0}))$ holds at all points $z_{0}\in\Omega$ satisfying (3.16) and (3.17). Because this includes $\mathscr{L}^{2}$ almost all points in $\Omega$ the proof of Step 1 is finished.
@@
$\int_{\mathbb{D}}\!|\textup{D}v_{j,r}|^{2}\,\mathrm{d}m(z)$ $=$ $\frac{\pi}{\rho_{r}^{2}}\mathchoice{{\vbox{\hbox{$\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\scriptstyle-$}}\kern-3.25pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.29166pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-1.875pt}}\!\int_{\mathbb{D}_{r}(z_{0})}\!|\textup{D}u_{j}|^{2}\,\mathrm{d}m(z)\leq\frac{\pi K}{\rho_{r}^{2}}\mathchoice{{\vbox{\hbox{$\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\scriptstyle-$}}\kern-3.25pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.29166pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-1.875pt}}\!\int_{\mathbb{D}_{r}(z_{0})}\!\det\textup{D}u_{j}\,\mathrm{d}m(z)$
$\to$ $\frac{\pi K}{\rho_{r}^{2}}\mathchoice{{\vbox{\hbox{$\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\scriptstyle-$}}\kern-3.25pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.29166pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-1.875pt}}\!\int_{\mathbb{D}_{r}(z_{0})}\!\det\textup{D}u\,\mathrm{d}m(z)\leq\frac{\pi K}{\rho_{r}^{2}}\mathchoice{{\vbox{\hbox{$\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\scriptstyle-$}}\kern-3.25pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.29166pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-1.875pt}}\!\int_{\mathbb{D}_{r}(z_{0})}\!|\textup{D}u|^{2}\,\mathrm{d}m(z)$
$\leq$ $\pi K\left(\rho_{r}^{-p}\mathchoice{{\vbox{\hbox{$\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\scriptstyle-$}}\kern-3.25pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.29166pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-1.875pt}}\!\int_{\mathbb{D}_{r}(z_{0})}\!|\textup{D}u|^{p}\,\mathrm{d}m(z)\right)^{\frac{2}{p}}$
@@
Now $v_{j}\colon\mathbb{D}\to\mathbb{C}$ are $K$-quasiregular maps and so we may Stoilow factorize as $v_{j}=h_{j}\circ f_{j}$, where $f_{j}\colon\mathbb{C}\to\mathbb{C}$ are $K$-quasiconformal principal and $h_{j}\colon f_{j}(\mathbb{D})\to\mathbb{C}$ are holomorphic. Taking a further subsequence if necessary (not relabelled) we may assume that $f_{j}\to f$ uniformly on $\mathbb{C}$, where $f\colon\mathbb{C}\to\mathbb{C}$ is a $K$-quasiconformal principal map. We then also have that $h_{j}\to 0$ locally uniformly on $f(\mathbb{D})$. It follows that $h_{j}^{\prime}\circ f_{j}\to 0$ uniformly on $\mathbb{D}_{s}(0)$ and therefore
@@
For the Dirichlet classes we have the following immediate corollary.
@@
Let $\Omega$ be a bounded open subset of $\mathbb{C}$ and $g\colon\mathbb{C}\to\mathbb{C}$ be a $K$-quasiregular map. If $u_{j}\in g+\mathrm{W}^{1,p_{K}}_{0}(\Omega)$ are $K$-quasiregular and $u_{j}\to u$ weakly in $\mathrm{W}^{1,p_{K}}(\Omega)$, then $u\in g+\mathrm{W}^{1,p_{K}}_{0}(\Omega)$ is $K$-quasiregular and
$\liminf_{j\to\infty}\int_{\Omega}\!\mathbf{B}_{p_{K}}(\textup{D}u_{j})\,\mathrm{d}m(z)\geq\int_{\Omega}\!\mathbf{B}_{p_{K}}(\textup{D}u)\,\mathrm{d}m(z).$

### v3

@@ 3.2. Lower semicontinuity with integrable distortion
where the last inequality follows by the assumed principal $\mathrm{W}^{1,1}$ quasiconvexity. Inserting the Jensen inequality for each $\nu_{z}$ in (3.14) yields the required lower semicontinuity. ∎
Proof of Corollary 1.7.
We start by establishing a coercivity inequality and note that the growth condition (1.17) implies that ${\bf E}(A)\geq-C\bigl{(}1+|A|^{p}+K_{A}+|\log J_{A}|\bigr{)}$ holds for all $A\in\mathbb{R}^{2\times 2}_{+}$. Consequently, invoking Young’s inequality in a routine manner we find positive constants $c_{1}$, $c_{2}>0$ such that ${\bf E}(A)+{\bf P}(A)\geq c_{1}|A|^{2}+c_{1}K_{A}^{q}-c_{2}-C|\log J_{A}|$ holds for all $A\in\mathbb{R}^{2\times 2}_{+}$. Next, combining the bound (3.13) with [5, Proposition 12.8] as in the proof of Proposition 3.5 above we arrive at
$\int_{\Omega}\!\bigl{(}{\bf E}(\textup{D}u)+{\bf P}(\textup{D}u)\bigr{)}\,\mathrm{d}m(z)\geq c_{0}\int_{\Omega}\!\bigl{(}|\textup{D}u|^{2}+K_{u}^{q}\bigr{)}\,\mathrm{d}m(z)-C_{0}$
for all $u\in\mathcal{A}_{g}$, where $c_{0}$, $C_{0}>0$ are positive constants that in particular are independent of $u$. Again using Young’s inequality in a standard manner we infer that the energy including the forcing term,
$\int_{\Omega}\!\bigl{(}{\bf E}(\textup{D}u)+{\bf P}(\textup{D}u)\bigr{)}\,\mathrm{d}m(z)+\bigl{\langle}F,u\bigr{\rangle}$
is bounded below on $\mathcal{A}_{g}$ and that, since ${\bf E}(\textup{D}g)+{\bf P}(\textup{D}g)$ is integrable over $\Omega$ by assumption, any minimizing sequence $(u_{j})\subset\mathcal{A}_{g}$ is bounded in $\mathrm{W}^{1,2}(\Omega)$ and has distortions $(K_{u_{j}})$ bounded in $\mathrm{L}^{q}(\Omega)$. By general principles we may then extract a subsequence (for convenience not relabelled) such that $u_{j}\to u$ weakly in $\mathrm{W}^{1,2}(\Omega)$. It then follows that $u\in g+\mathrm{W}^{1,2}_{0}(\Omega)$ and, see [8, Ch. 21], that $u\colon\Omega\to g(\Omega)$ is a homeomorphism with $K_{u}\in\mathrm{L}^{q}(\Omega)$. In particular, $u\in\mathcal{A}_{g}$ and by [15] and Theorem 3.4 $u$ is therefore a minimizer, as required. ∎
@@ 3.3.1. An integrability property of the Burkholder functional
(3.16) $\fint_{\mathbb{D}_{\delta R}(z_{0})}\!\mathbf{B}_{p}(\textup{D}u)\,\mathrm{d}m(z)\geq-\frac{C}{R^{pK}(1-\delta)^{pK}}\sup_{\mathbb{D}_{R}(z_{0})}|u|^{p}.$
@@
Changing variables we may without loss of generality assume that $z_{0}=0$ and $R=1$. We then Stoilow factorize $u$ on $\mathbb{D}$, whereby $u=h\circ f$ for a $K$-quasiconformal normalized principal map $f\colon\mathbb{C}\to\mathbb{C}$ and a holomorphic map $h\colon f(\mathbb{D})\to\mathbb{C}$. If $m\equiv\sup_{\mathbb{D}}|u|$, then also $m=\sup_{f(\mathbb{D})}|h|$ and so by Cauchy’s integral formula and Hölder properties of $K$-quasiconformal principal maps [8] we estimate
@@
We split the proof into two steps and remark that the assumption $u\in\mathrm{W}^{1,p}_{\operatorname{loc}}(\Omega)$ only is used in the second step.
[…]
@@
Step 1. $\frac{\mathrm{d}\lambda}{\mathrm{d}\mathscr{L}^{2}}\geq\mathbf{B}_{p}(\textup{D}u)$ holds $\mathscr{L}^{2}$ almost everywhere in $\Omega$. The proof consist of a rather general localization argument at a point of differentiabilty of $u$. Arguments of this sort are known in the literature as blow-up arguments and in the present context go back to I. Fonseca and S. Müller [40]. The fact that the Burkholder functional is nonpositive on its effective domain means that one must be more careful than in the classical set-up, where the considered functionals are bounded below and therefore concentration effects are not so harmful.
@@
[…]
@@
$\lim_{E\notowner r\searrow 0}\lim_{j\to\infty}\fint_{\mathbb{D}_{r}(z_{0})}\!\mathbf{B}_{p}(\textup{D}u_{j})\,\mathrm{d}m(z)=-\frac{\mathrm{d}\lambda}{\mathrm{d}\mathscr{L}^{2}}(z_{0}),$
@@
Finally, taking $s\nearrow 1$ we conclude that $\frac{\mathrm{d}\lambda}{\mathrm{d}\mathscr{L}^{2}}(z_{0})\geq\mathbf{B}_{p}(\textup{D}u(z_{0}))$ holds at all points $z_{0}\in\Omega$ satisfying (3.17) and (3.18). Because this includes $\mathscr{L}^{2}$ almost all points in $\Omega$ the proof of Step 1 is finished.
@@
$\int_{\mathbb{D}}\!|\textup{D}v_{j,r}|^{2}\,\mathrm{d}m(z)$ $=$ $\frac{\pi}{\rho_{r}^{2}}\fint_{\mathbb{D}_{r}(z_{0})}\!|\textup{D}u_{j}|^{2}\,\mathrm{d}m(z)\leq\frac{\pi K}{\rho_{r}^{2}}\fint_{\mathbb{D}_{r}(z_{0})}\!\det\textup{D}u_{j}\,\mathrm{d}m(z)$
$\to$ $\frac{\pi K}{\rho_{r}^{2}}\fint_{\mathbb{D}_{r}(z_{0})}\!\det\textup{D}u\,\mathrm{d}m(z)\leq\frac{\pi K}{\rho_{r}^{2}}\fint_{\mathbb{D}_{r}(z_{0})}\!|\textup{D}u|^{2}\,\mathrm{d}m(z)$
$\leq$ $\pi K\left(\rho_{r}^{-p}\fint_{\mathbb{D}_{r}(z_{0})}\!|\textup{D}u|^{p}\,\mathrm{d}m(z)\right)^{\frac{2}{p}}$
@@
Now $v_{j}\colon\mathbb{D}\to\mathbb{C}$ are $K$-quasiregular maps and so we may Stoilow factorize as $v_{j}=h_{j}\circ f_{j}$, where $f_{j}\colon\mathbb{C}\to\mathbb{C}$ are $K$-quasiconformal normalized principal and $h_{j}\colon f_{j}(\mathbb{D})\to\mathbb{C}$ are holomorphic. Taking a further subsequence if necessary (not relabelled) we may assume that $f_{j}\to f$ uniformly on $\mathbb{C}$, where $f\colon\mathbb{C}\to\mathbb{C}$ is a $K$-quasiconformal normalized principal map. We then also have that $h_{j}\to 0$ locally uniformly on $f(\mathbb{D})$. It follows that $h_{j}^{\prime}\circ f_{j}\to 0$ uniformly on $\mathbb{D}_{s}(0)$ and therefore
@@
For the Dirichlet classes we have the following almost immediate corollary.
@@
Let $p>2$ and put $K\equiv p/(p-2)$. Let $\Omega$ be a bounded open subset of $\mathbb{C}$ and $g\colon\mathbb{C}\to\mathbb{C}$ be a $K$-quasiregular map of class $\mathrm{W}^{1,p}_{\operatorname{loc}}(\mathbb{C})$ If $u_{j}\in g+\mathrm{W}^{1,p}_{0}(\Omega)$ are $K$-quasiregular and $u_{j}\to u$ weakly in $\mathrm{W}^{1,p}(\Omega)$, then $u\in g+\mathrm{W}^{1,p}_{0}(\Omega)$ is $K$-quasiregular and
$\liminf_{j\to\infty}\int_{\Omega}\!\mathbf{B}_{p}(\textup{D}u_{j})\,\mathrm{d}m(z)\geq\int_{\Omega}\!\mathbf{B}_{p}(\textup{D}u)\,\mathrm{d}m(z).$
We emphasize that we assume $u_{j}\to u$ weakly in the critical Sobolev space $\mathrm{W}^{1,p}(\Omega)$ and that this of course automatically entails that $u\in\mathrm{W}^{1,p}(\Omega)$. If we only knew that $_{j}, uu$ were $K$-quasiregular and $u_{j}\to u$ uniformly on $\Omega$, then we would merely have that $\textup{D}u$ belonged to the local Marcinkiewicz space $\mathrm{L}^{p,\infty}_{\operatorname{loc}}(\Omega)$. Theorem 1.5 does not cover this situation.
Proof.
Consider the metric neighbourhood $\Omega_{r}\equiv\mathbb{D}_{r}(\Omega)\equiv\bigl{\{}z\in\mathbb{C}:\,\mathrm{dist}(z,\Omega)<r\bigr{\}}$ for $r>0$. It is clearly precompact and $\Omega\Subset\Omega_{r}$. Since the function $\mathrm{dist}(\cdot,\Omega)$ is Lipschitz we may use the coarea formula to see that also $\mathscr{L}^{2}(\partial\Omega_{r})=0$ holds for $\mathscr{L}^{1}$ almost all $r>0$, and we fix such an $r$. Now extend the maps $_{j}, uu$ to $\mathbb{C}\setminus\Omega$ by $g$, and record that hereby $_{j}, uu$ are $K$-quasiregular maps in the critical Sobolev space $\mathrm{W}^{1,p}_{\operatorname{loc}}(\mathbb{C})$ and that $u_{j}\to u$ weakly in $\mathrm{W}^{1,p}_{\operatorname{loc}}(\mathbb{C})$. The latter in particular implies that also $u_{j}\to u$ uniformly on $\mathbb{C}$, and so according to Theorem 1.5
$\liminf_{j\to\infty}\int_{\Omega_{r}}\!\mathbf{B}_{p}(\textup{D}u_{j})\,\mathrm{d}m(z)\geq\int_{\Omega_{r}}\!\mathbf{B}_{p}(\textup{D}u)\,\mathrm{d}m(z)$
holds. Note that here we have $u_{j}=u=g$ on $\Omega_{r}\setminus\Omega$ and so by standard properties of Sobolev functions and approximate derivatives we get in particular that $\textup{D}u_{j}=\textup{D}u$ a.e. on $\Omega_{r}\setminus\Omega$ (also in the pathological case where $\mathscr{L}^{2}(\partial\Omega)>0$ that has not been excluded). Because $\mathbf{B}_{p}(\textup{D}u)\in\mathrm{L}^{1}_{\operatorname{loc}}(\mathbb{C})$ the required conclusion follows from this. ∎
Acknowledgments